OpenStudy (anonymous):

Given the points (-2, 1) and (5, -5): a) Find the slope of the line through the points. b) Write an equation in point-slope form of the line through the points. c) Convert the equation to slope-intercept form. d) Convert the equation to standard form. e) Graph the equation. You may use the axes provided, or create your own graph

OpenStudy (anonymous):

I'll help, but I won't do all the work. Slope: \[m =\frac{ y2 - y1 }{ x2 - x1 } \to \frac{ -5 - 1 }{ 5 - (-2) } \to \frac{ -5 -1 }{ 5 + 2 }\] What's the slope?

OpenStudy (anonymous):

*I meant: What is (-5) - 1? What is 5 + 2?

OpenStudy (anonymous):

@jmays14